how to calculate the neet score?
suppose i attempted 100 questions and among them only 50 are correct wht will be my score?
 
 
 

Grade:12th pass

1 Answers

Umakant Biswal
6 years ago
dear student 
its very simple process 
acc to the instructions , you will get +4 marks for each correct answer and -1 for each wrong answer and 0 for unattempted ques . 
so, you attempted 100 ques and 50 are correct means 50 are incorrect . 
so, correct answer = 50* 4 = 200 marks . 
and 50 incorrect answer . = 200-50 = 150 
so, your score will be 150 here . 
HOPE IT CLEARS 
regards 
kindly approve